# Losing Visual Needles in Image Haystacks: Vision Language Models are Easily Distracted in Short and Long Contexts

### v1

@@ Abstract
Across these tasks, a diverse set of VLMs rapidly lose performance as the visual context length grows, often exhibiting a striking exponential decay trend. This test assesses how well VLMs can ignore irrelevant information when answering queries—a task that is quite easy for language models (LMs) in the text domain—demonstrating that current state-of-the-art VLMs lack this essential capability for many long-context applications.
@@
Figure 1: The impact of visual context on vision-language models (VLMs) in our modified, multi-image versions of the OK-VQA, MMStar, and MMBench evaluation benchmarks. Distractor images around the target image increase the visual context length needed to answer the questions. VLM performance exhibits an exponential decay against distractor count, evident in both single composed (cmp) and multiple interleaved (int) input image configurations.
@@ 4 Experiments
Figure 4: VLM Performance on MMStar and OK-VQA. Note the clearly declining exponential fit trends for many of the models. The (model, task) pairs for which these trends do not hold by and large are below the random baseline.
@@ 5 Results
Across all models on OK-VQA and the majority on MMStar subsets for SeedBench, MMBench, ScienceQA, and AI2D, we observe striking exponential decay trends in model performance with increasing visual context length. Correlation coefficients, $r^{2}$, and $p$-values for each trendline are reported in Table 3, Appendix D. In general, the closed-source models outperform their open-weight counterparts, especially on multiple-choice tasks. Among the open-weight models, PaliGemma performs the best, likely due to its substantially more extensive training on vision-language tasks compared to its counterparts like LLava or Mantis.
@@
The most noteworthy point is this: the exponential decay trend holds equally well in interleaved and composed settings. This indicates that the performance-visual context length trend is fundamental and cannot be attributed solely to downsampling effects in the composite setting. Furthermore, for the models tested in both conditions (GPT-4V and Gemini), the same trends are observed in both interleaved and composite settings. Sub-chance performance on the multiple-choice question-answering datasets, particularly for the closed-weight models, occur due to high refusal rates. This may illustrate how “alignment hampering” can hinder performance.
@@ 7 Discussion
Why do we observe such striking exponential decay in performance with increasing visual context length? The fact that the trend is consistent in both interleaved and composite settings that downsampling effects in the latter are not the primary cause of performance decay. There may be an information-theoretic interpretation of this behavior: an increasing signal-to-noise ratio. As the visual context length increases, the amount of signal (relevant information) remains constant, while the amount of noise (distractor information) increases.
@@ 8 Conclusion
Although LoCoVQA is a generalized process applicable to any VLM benchmark, we only evaluated it on three tasks. While the strong exponential decay trends between visual context length and performance observed across all three are compelling, expansion of LoCoVQA to additional tasks would render the findings clearer.
@@ Appendix E Supplementary Results
Table 2: Exponential curve fit $r^{2}$ values are reported for each open-weight model, followed by a symbol denoting the $p$-value for statistical significance. The symbols represent:  for $p\leq 0.05$, ∗∗ for $p\leq 0.01$, & ∗∗∗ for $p\leq 0.001$. Light pink cells represent correlations for sets of values that fall below chance performance.
@@
Table 3: Exponential curve fit $r^{2}$ values are reported for each closed-source model, followed by a symbol denoting the $p$-value for statistical significance. The symbols represent:  for $p\leq 0.05$, ∗∗ for $p\leq 0.01$, & ∗∗∗ for $p\leq 0.001$. Light pink cells represent correlations for sets of values that fall below chance performance.

### v2

@@ Abstract
Across these tasks, a diverse set of VLMs rapidly lose performance as the visual context length grows, often exhibiting a striking logarithmic decay trend. This test assesses how well VLMs can ignore irrelevant information when answering queries—a task that is quite easy for language models (LMs) in the text domain—demonstrating that current state-of-the-art VLMs lack this essential capability for many long-context applications.
@@
Figure 1: The impact of visual context on vision-language models (VLMs) in our modified, multi-image versions of the OK-VQA, MMStar, and MMBench evaluation benchmarks. Distractor images around the target image increase the visual context length needed to answer the questions. VLM performance exhibits a logarithmic decay against distractor count, evident in both single composed (cmp) and multiple interleaved (int) input image configurations.
@@ 4 Experiments
Figure 4: VLM Performance on MMStar and OK-VQA. Note the clearly declining logarithmic fit trends for many of the models. The (model, task) pairs for which these trends do not hold by and large are below the random baseline.
@@ 5 Results
Across all models on OK-VQA and the majority on MMStar subsets for SeedBench, MMBench, ScienceQA, and AI2D, we observe striking logarithmic decay trends in model performance with increasing visual context length. Correlation coefficients, $r^{2}$, and $p$-values for each trendline are reported in Table 3, Appendix D. In general, the closed-source models outperform their open-weight counterparts, especially on multiple-choice tasks. Among the open-weight models, PaliGemma performs the best, likely due to its substantially more extensive training on vision-language tasks compared to its counterparts like LLava or Mantis.
@@
The most noteworthy point is this: the logarithmic decay trend holds equally well in interleaved and composed settings. This indicates that the performance-visual context length trend is fundamental and cannot be attributed solely to downsampling effects in the composite setting. Furthermore, for the models tested in both conditions (GPT-4V and Gemini), the same trends are observed in both interleaved and composite settings. Sub-chance performance on the multiple-choice question-answering datasets, particularly for the closed-weight models, occur due to high refusal rates. This may illustrate how “alignment hampering” can hinder performance.
@@ 7 Discussion
Why do we observe such striking logarithmic decay in performance with increasing visual context length? The fact that the trend is consistent in both interleaved and composite settings that downsampling effects in the latter are not the primary cause of performance decay. There may be an information-theoretic interpretation of this behavior: an increasing signal-to-noise ratio. As the visual context length increases, the amount of signal (relevant information) remains constant, while the amount of noise (distractor information) increases.
@@ 8 Conclusion
Although LoCoVQA is a generalized process applicable to any VLM benchmark, we only evaluated it on three tasks. While the strong logarithmic decay trends between visual context length and performance observed across all three are compelling, expansion of LoCoVQA to additional tasks would render the findings clearer.
@@ Appendix E Supplementary Results
Table 2: Logarithmic curve fit $r^{2}$ values are reported for each open-weight model, followed by a symbol denoting the $p$-value for statistical significance. The symbols represent:  for $p\leq 0.05$, ∗∗ for $p\leq 0.01$, & ∗∗∗ for $p\leq 0.001$. Light pink cells represent correlations for sets of values that fall below chance performance.
@@
Table 3: Logarithmic curve fit $r^{2}$ values are reported for each closed-source model, followed by a symbol denoting the $p$-value for statistical significance. The symbols represent:  for $p\leq 0.05$, ∗∗ for $p\leq 0.01$, & ∗∗∗ for $p\leq 0.001$. Light pink cells represent correlations for sets of values that fall below chance performance.